How many ways are there to select 15 countries in the United Nations to serve on a council if 4 are selected from a block of 35, 5 are selected from a block of 40, and the others are selected from the remaining 55 countries?




1
Expert's answer
2022-01-30T16:24:03-0500

The order does not matter


"N=\dbinom{35}{4}\dbinom{40}{5}\dbinom{55}{15-4-5}"

"=\dfrac{35!}{4!(35-4)!}\cdot \dfrac{40!}{5!(40-5)!}\cdot \dfrac{55!}{6!(55-6)!}"

"=52360\cdot 658008\cdot28989675"

"=998,789,937,209,064,000(ways)"